# OpenGait: A Comprehensive Benchmark Study for Gait Recognition towards Better Practicality

### v1

@@ VI-A Comparison Around DeepGaitV2
Except for the degradation observed on CASIA-B [10] (Table VII), DeepGaitV2 outperforms other silhouette-based SoTA methods across all other datasets, as demonstrated in Table VIII, X, and XI. Particularly noteworthy is its performance on real-world large-scale gait datasets such as Gait3D [9] and GREW [8], where DeepGaitV2 exhibits more substantial improvements compared to its performance on constrained datasets like OU-MVLP [11], SUSTech1K [21], and CCPG [20]. This alignment with the primary objective of our benchmark study underscores its role in advancing gait recognition toward enhanced practicality.
@@
Moreover, Table VII illustrates that the performance advantage of DeepGaitV2 primarily arises from its depth, i.e., the larger size, and the higher rank-1 accuracy, before reaching the marginal effect. While this observation may seem intuitive for vision tasks, it has been often overlooked in previous gait studies, as discussed in Sec. V-B. Therefore, this work emphasizes the importance of adopting powerful network architectures as the backbone for future deep gait model designs.
VI-B Comparison Around SkeletonGait
@@
As shown in Table VIII, X, and XI, SkeletonGait outperforms the latest skeleton-based methods by breakthrough improvements in most cases. It gains $+5.3\%$, $+15.6\%$, $+22.9\%$, $29.2\%$, and $+17.4\%$ rank-1 accuracy on the OU-MVLP [11], Gait3D [9], GREW [8], SUSTech1K [21], and CCPG [20] datasets, respectively.
@@
To exclude the potential positive influence brought by the model size of SkeletonGait, we reduce its channels by half, thus making its model size nearly identical to that of GPGait [29], i.e., 2.85 v.s. 2.78M. After that, SkeletonGait reached the rank-1 accuracy of 33.2% and 70.9% on Gait3D [9] and GREW [8], still maintaining a higher result than prior skeleton-based methods shown in Fig. VIII.
@@
To demonstrate the robustness of SkeletonGait to different upstream pose estimators, we conduct experiments using both the officially provided AlphaPose and OpenPose data on OU-MVLP [11], resulting in a rank-1 accuracy of 67.4% and 65.9%, respectively. These two results consistently surpass other SoTA skeleton-based methods, revealing the robustness of SkeletonGait. Table VIII presents the higher one.
@@
Input Method Venue Probe Sequence (R-1) Overall
@@
Silhouette- based GaitSet [7] AAAI’19 69.1 68.2 37.4 65.0 63.1 61.0 67.2 23.0 65.0 84.8
GaitPart [15] CVPR’19 62.2 62.8 33.1 59.5 57.2 54.8 57.2 21.7 59.2 80.8
GaitGL [16] ICCV’21 67.1 66.2 35.9 63.3 61.6 58.1 66.6 17.9 63.1 82.8
GaitBase [17] CVPR’23 81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
@@
Skeleton- based GaitGraph2 [30] CVPRW’22 22.2 18.2 6.8 18.6 13.4 19.2 27.3 16.4 18.6 40.2
Gait-TR [64] ES’23 33.3 31.5 21.0 30.4 22.7 34.6 44.9 23.5 30.8 56.0
MSGG [35] MTA’23 67.11 66.16 35.92 63.31 61.58 58.07 66.59 17.88 33.8 -
@@
Multi- modal BiFusion [35] MTA’23 69.8 62.3 45.4 60.9 54.3 63.5 77.8 33.7 62.1 83.4
ParsingGait [25] MM’23 82.7 73.3 30.4 76.4 70.8 74.6 84.3 50.5 75.0 90.7
@@
Input Model Venue Gait Evaluation Protocol ReID Evaluation Protocol
@@
Silhouette- based GaitSet [7] AAAI’19 60.2 65.2 65.1 68.5 64.8 77.5 85.0 82.9 87.5 83.2
GaitPart [15] CVPR’20 64.3 67.8 68.6 71.7 68.1 79.2 85.3 86.5 88.0 84.8
AUG-OGBase [20] CVPR’23 52.1 57.3 60.1 63.3 58.2 70.2 76.9 80.4 83.4 77.7
GaitBase [17] CVPR’23 71.6 75.0 76.8 78.6 75.5 88.5 92.7 93.4 93.2 92.0
@@
Skeleton- based GaitGraph2 [30] CVPRW’22 5.0 5.3 5.8 6.2 5.1 5.0 5.7 7.3 8.8 6.7
Gait-TR [64] ES’23 15.7 18.3 18.5 17.5 17.5 24.3 28.7 31.1 28.1 28.1
MSGG [35] MTA’23 29.0 34.5 37.1 33.3 33.5 43.1 52.9 57.4 49.9 50.8
@@
Multi- modal BiFusion [35] MTA’23 62.6 67.6 66.3 66.0 65.6 77.5 84.8 84.8 82.9 82.5
ParsingGait [25] MM’23 55.3 58.9 64.0 66.7 61.2 73.5 78.4 85.2 87.0 81.0
@@
Figure 9: The rank-1 accuracy of SkeletonGait on OU-MVLP [11] over the probe-gallery view pairs.
@@
As discussed in Sec. V-C, the skeleton map can be perceived as a silhouette devoid of body shape information. Through a detailed comparison of the performance between SkeletonGait and DeepGaitV2, we obtain the following thoughts:
@@
  • Importance of Structural Features. Structural features play a more important role than previously shown. They contribute to over 50% of the overall performance, as indicated by the rank-1 accuracy ratios between SkeletonGait and DeepGaitV2.
@@
  • Superiority of Skeleton Data. In situations where silhouette data might be less reliable, such as the night case of SUSTech1K in Table X, SkeletonGait exhibits a considerable performance advantage over DeepGaitV2, highlighting the potential of skeleton data in such cases.
@@
  • Cross-view Challenge. The results on the diagonal and anti-diagonal of Fig. 9 respectively present the performance of SkeletonGait in identical- and nearly symmetric-view cases, where other regions reflect cross-view conditions. As we can see, the cross-view scenarios remain a major challenge for SkeletonGait.
@@
  • Concerns on GREW. The GREW dataset is widely acknowledged as the most challenging gait dataset due to its largest scale and real-world settings. However, SkeletonGait achieves a comparable performance compared to DeepGaitV2 on GREW, rather than on other datasets, as shown in Table VIII. Considering SkeletonGait works well on the cross-limited-view cases as shown in Fig. 9, we assume that GREW may contain not enough cross-view covariates. Further investigations are warranted.
@@
The highest performance is in bold.
VI-C Comparison Around SkeletonGait++
As demonstrated in Table VIII, X, and XI, SkeletonGait++ achieves a new SoTA with notable improvements compared to other multi-modal methods. Specifically, it gains $+1.4\%$, $+21.2\%$, $+9.5\%$, and $23.2\%$ rank-1 accuracy on the Gait3D [9], GREW [8], SUSTech1K [21], and CCPG [20] datasets, respectively. Here the result missing on OU-MVLP [11] is due to the lack of frame-by-frame alignment of skeleton and silhouette inputs for this dataset.
To better understand the differences among DeepGaitV2, SkeletonGait, and SkeletonGait++, Fig. 10 visualizes their activation maps through the CAM technique [68]. As a result, DeepGaitV2 pays more attention to regions that exhibit distinct and discriminative body shapes. On the other hand, SkeletonGait can only concentrate on ‘clean’ structural features over the body joints and limbs. In comparison, SkeletonGait++ strikes a balance between these approaches, effectively capturing the ‘comprehensive’ gait patterns that are rich in both body shape and structural characteristics. Especially for the challenging cases shown in Fig. 10 (b), SkeletonGait++ adaptively leverages the skeleton branch to support the robust gait representation learning. This is an urgent need for practical applications, and we also think this is the main reason causing the SkeletonGait++’s performance gains on Gait3D [9] and GREW [8] datasets.
Figure 10: The activation visualization of DeepGaitV2 vs. SkeletonGait and SkeletonGait++ on SUSTech1K [21], Gait3D [9], and GREW [8].
VI-D Ablation Study
The preceding paragraphs have thoroughly explored the relationship and distinction among DeepGaitV2, SkeletonGait, and SkeletonGait++, through comprehensive experiments across all the utilized datasets. As a supplementary analysis, here we further investigate their robustness to the hyper-parameters as outlined below.
To determine the suitable network capacity for DeepGaitV2, Table VII has deepened the backbone depth from 10 to 30 layers. We further widen the network, and the results in Table XII (a) exhibit limited performance improvements but introduce significant training parameters and computation costs. Therefore, we set the network width $C$ to 64 as default for a better accuracy-computation trade-off.
To determine the empirical value of $\sigma$ in generating the skeleton map as described in Eq. (4) and (5), we range it from 4 to 16, and Table XII (b) shows that: a) SkeletonGait is robust to the value of $\sigma$. b) $\sigma=8.0$ is an experimentally optimal choice. These findings also underscore the robustness and discriminative capacity of the proposed skeleton map as a gait modality used for human identification.
To determine the fusion location and mode within SkeletonGait++ shown in Fig. 6, we try all the combinations, and Table XII (c) reveals that: a) SkeletonGait++ is robust to both fusion location and mode. b) The low-level attention fusion is an experimentally optimal choice. Moreover, these observations also highlight the promising potential of multi-modal gait recognition methods.
VII Conclusion and Future Work

### v2

@@ VI-A Analysis by the Appearance-based DeepGaitV2
Except for the degradation observed on CASIA-B [10] (Table VII), DeepGaitV2 outperforms other silhouette-based methods across all other datasets, as demonstrated in Table IX, X, and XI. Particularly, DeepGaitV2 exhibits more substantial improvements on real-world large-scale gait datasets, such as Gait3D [9] and GREW [8], than on constrained datasets like OU-MVLP [11], SUSTech1K [21], and CCPG [20]. This alignment with the primary objective of our benchmark study underscores its role in advancing gait recognition in real-world applications.
@@
Moreover, Table VII illustrates that the performance advantage of DeepGaitV2 primarily arises from its depth. The deeper and the larger, the higher rank-1 accuracy can be gained. It is straightforward in deep learning for computer vision but was not well addressed in gait recognition. The overfitting on some small constrained datasets like CASIA-B can mislead the exploring to a deep model. Therefore, this work emphasizes the importance of adopting powerful network architectures as the backbone in gait recognition.
VI-B Analysis by the Model-based SkeletonGait
@@
As shown in Table IX, X, and XI, SkeletonGait outperforms the latest skeleton-based methods by breakthrough improvements in most cases. It gains $+5.3\%$, $+15.6\%$, $+22.9\%$, $29.2\%$, and $+17.4\%$ rank-1 accuracy on OU-MVLP [11], Gait3D [9], GREW [8], SUSTech1K [21] and CCPG [20], respectively.
@@
To exclude the potential positive influence brought by the model size of SkeletonGait, we reduced its channels by half. It made its model size nearly identical to that of GPGait [34], i.e., 2.85M v.s. 2.78M. Even so, SkeletonGait can still reach the rank-1 accuracy of 33.2% on Gait3D [9] and 70.9% on GREW [8], and achieves better results than other skeleton-based methods shown in Table IX.
@@
To demonstrate the robustness of SkeletonGait to different upstream pose estimators, we conduct experiments using both AlphaPose and OpenPose data provided by OU-MVLP [11], resulting in a rank-1 accuracy of 67.4% and 65.9%, respectively. These two results consistently surpass other skeleton-based methods, revealing the robustness of SkeletonGait. Note Table IX only presents the higher one for brevity.
@@
Input Method Year Probe Sequence (R-1) Overall
@@
Silhouette- based GaitSet [7] 2019 69.1 68.2 37.4 65.0 63.1 61.0 67.2 23.0 65.0 84.8
GaitPart [15] 2019 62.2 62.8 33.1 59.5 57.2 54.8 57.2 21.7 59.2 80.8
GaitGL [16] 2021 67.1 66.2 35.9 63.3 61.6 58.1 66.6 17.9 63.1 82.8
GaitBase [17] 2023 81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
@@
Skeleton- based GaitGraph2 [35] 2022 22.2 18.2 6.8 18.6 13.4 19.2 27.3 16.4 18.6 40.2
Gait-TR [71] 2023 33.3 31.5 21.0 30.4 22.7 34.6 44.9 23.5 30.8 56.0
MSGG [42] 2023 67.1 66.2 35.9 63.3 61.6 58.1 66.6 17.9 33.8 -
@@
Multi- modal BiFusion [42] 2023 69.8 62.3 45.4 60.9 54.3 63.5 77.8 33.7 62.1 83.4
@@
Input Model Year Gait Evaluation Protocol ReID Evaluation Protocol
@@
Silhouette- based GaitSet [7] 2019 60.2 65.2 65.1 68.5 64.8 77.5 85.0 82.9 87.5 83.2
GaitPart [15] 2020 64.3 67.8 68.6 71.7 68.1 79.2 85.3 86.5 88.0 84.8
AUG-OGBase [20] 2023 - - 84.7 88.4 89.4 - -
GaitBase [17] 2023 71.6 75.0 76.8 78.6 75.5 88.5 92.7 93.4 93.2 92.0
@@
Skeleton- based GaitGraph2 [35] 2022 5.0 5.3 5.8 6.2 5.1 5.0 5.7 7.3 8.8 6.7
Gait-TR [71] 2023 15.7 18.3 18.5 17.5 17.5 24.3 28.7 31.1 28.1 28.1
MSGG [42] 2023 29.0 34.5 37.1 33.3 33.5 43.1 52.9 57.4 49.9 50.8
@@
Multi- modal BiFusion [42] 2023 62.6 67.6 66.3 66.0 65.6 77.5 84.8 84.8 82.9 82.5
ParsingGait [30] 2023 55.3 58.9 64.0 66.7 61.2 73.5 78.4 85.2 87.0 81.0
@@
Note: this table presents the rank-1 accuracy.
@@
Figure 9: The accuracies of SkeletonGait over all probe-gallery view pairs on OU-MVLP. The horizontal and vertical coordinates respectively represent the filming viewpoints of the probe and gallery set.
@@
As discussed in the previous section, the skeleton map can be perceived as a silhouette devoid of body shape. By comparing the rank-1 accuracy of DeepGaitV2 vs. SkeletonGait in Table IX, X, and XI, i.e., 91.9% vs. 67.4% on OU-MVLP [11], 74.4% vs. 38.1% on Gait3D [9], 77.7% vs. 77.4% on GREW [8], 82.3% vs. 63.0% SUSTech1K [21], and 83.3% vs. 50.9% on CCPG [20], we obtain the following thoughts:
@@
  • Importance of Structural Features. Structural features play an important role. They contribute to over 50% of the overall performance, as indicated by the rank-1 accuracy ratios between SkeletonGait and DeepGaitV2.
@@
  • Superiority of Skeleton Data. In situations where silhouettes are noisy, such as the data at night in SUSTech1K, SkeletonGait exhibits a considerable performance advantage over DeepGaitV2 (50.0% vs. 28.8%, Table X). It shows the robustness of the skeleton data in such cases.
@@
  • Cross-view Challenge. The results on the diagonal and anti-diagonal of Fig. 9 respectively are the rank-1 accuracies of SkeletonGait in identical- and nearly symmetric-view cases, where others reflect the cross-view situations on OU-MVLP [11]. As shown in the figure, the cross-view issue remains a major challenge for SkeletonGait.
@@
  • Concerns on GREW. SkeletonGait cannot get comparable results as DeepGaitV2 on OU-MVLP and Gait3D. It just achieves 67.4% (vs. 91.9% by DeepGaitV2) on OU-MVLP, and 38.1% (vs. 74.4% by DeepGaitV2) on Gait3D according to Table IX. But surprisingly it can achieve a comparable accuracy of 77.4% (vs. 77.7% by DeepGaitV2) on GREW. GREW has been widely acknowledged as the most challenging gait dataset due to its largest scale and real-world settings. Given that SkeletonGait performs well on the same view but poorly on cross views, as shown in the cross-view experiments on OU-MVLP (Fig. 9), one possibility is that GREW does not contain enough cross-view variations. However, further proofs are needed to rigorously confirm this point.
VI-C Analysis by the Multi-modal SkeletonGait++
As demonstrated in Table IX, X, and XI, SkeletonGait++ achieves a new SoTA with notable improvements compared to other single-modal and multi-modal methods. Specifically, it gains $+1.4\%$, $+21.2\%$, $+9.5\%$, and $23.2\%$ rank-1 accuracy on the Gait3D [9], GREW [8], SUSTech1K [21], and CCPG [20] datasets, respectively. The result missing on OU-MVLP in Table IX is caused by the lack of frame-by-frame alignment between the skeleton and silhouette, which was not provided by the dataset owner and cannot be obtained by some methods.
To better understand the differences among DeepGaitV2, SkeletonGait, and SkeletonGait++, Fig. 10 visualizes their activation maps through the CAM technique [74]. As a result, DeepGaitV2 pays more attention to regions that exhibit distinct and discriminative body shapes. On the other hand, SkeletonGait can only concentrate on clean structural features over the body joints and limbs. In comparison, SkeletonGait++ strikes a balance between these approaches, effectively capturing the comprehensive gait patterns that are rich in both body shape and structural characteristics. Especially for the challenging cases shown in Fig. 10 (b), SkeletonGait++ adaptively leverages the skeleton branch to support the robust gait representation learning. This is an urgent need for real applications. We also think it is the main reason for the improvement by SkeletonGait++ on Gait3D [9] and GREW [8].
Figure 10: The activation visualization of DeepGaitV2 vs. SkeletonGait and SkeletonGait++ (vertical axis) on some typical samples from SUSTech1K [21], Gait3D [9], and GREW [8] (horizontal axis).
VI-D Ablation Study on Hyper-parameters
The previous part of the section has thoroughly explored the relationship and distinction among DeepGaitV2, SkeletonGait, and SkeletonGait++, through comprehensive experiments across multiple popular gait datasets. The robustness of those methods to their hyper-parameters is analyzed below.
To determine a suitable network size for DeepGaitV2, the results by different depths (10, 14, 22, and 30 layers) are shown in Table VII. We also changed the number of channels from 32 to 128, and the results are in Table XII (a). Widen the network from 64 to 128 gains limited accuracy improvement but introduces significant training parameters and computation costs. Therefore, we set the number of channels $C$ to 64 for a better accuracy-computation trade-off.
To determine the empirical value of $\sigma$ in generating the skeleton map as described in Eq. (4) and (5), we tested it from 4 to 16, and the results in Table XII (b) shows that: i) SkeletonGait is robust to the value of $\sigma$. ii) $\sigma=8.0$ is an experimentally optimal choice. These findings also underscore the robustness and discriminative capacity of the proposed skeleton map as a gait modality for gait recognition.
To determine the fusion location and mode within SkeletonGait++ shown in Fig. 6, we tried all the combinations. The results in Table XII (c) reveals that: i) SkeletonGait++ is robust to both fusion locations and fusion modes. ii) The low-level attention fusion is an experimentally optimal choice. Moreover, these observations also highlight the promising potential of multi-modal gait recognition methods.
@@ VI-D Ablation Study on Hyper-parameters
Note: the highest performance is in bold.
VII Trends and Challenges
Despite significant efforts, we acknowledge that OpenGait has yet to address several crucial topics essential for advancing gait recognition toward broader real-world applicability. To bridge this gap, this section offers key insights into the current state of gait recognition and outlines a clear roadmap for future advancements, highlighting emerging trends and opportunities. Building on recent literature and feedback from the OpenGait community, we present the following discussion:
$\bullet$ Handling Noisy Inputs. Most widely used gait input modalities, such as silhouettes and skeletons, are typically extracted from raw videos using third-party upstream algorithms, including pedestrian segmentation and pose estimation. However, accurately extracting these modalities from less-than-ideal environments pose their own set of challenges. Current approaches to mitigate this issue fall into three main directions: (a) enhancing upstream models—though this falls outside the core scope of gait recognition research, (b) exploring more robust modalities, such as point clouds [21], (c) developing quality-robust gait models [75, 70], (d) leveraging the complementarity of multi-modal inputs to enhance gait representation [39], and (e) facilitating end-to-end modeling [29, 32, 76, 31, 77] to eliminate the inaccuracy accumulation. However, a comprehensive evaluation comparing the effectiveness of these strategies remains lacking, leaving an open question in the field.
$\bullet$ Ideal Gait Input. While RGB data can introduce distractions caused by variations in clothing, lighting, and backgrounds, they could also provide rich identity cues such as fine-grained body segmentation and 3D body priors. To enhance gait representations, recent studies have explored: (a) advanced sensors such as LiDAR for point cloud-based gait analysis [21, 78], (b) improved gait modalities, including human parsing [30, 23] and SMPL models [32, 9], (c) multi-modal frameworks integrating complementary gait representations [9, 79, 19, 80], and (d) end-to-end training strategies with inductive biases, such as appearance reconstruction [32, 76], differentiable silhouette edge modeling [31], and spatial feature smoothing [77]. OpenGait recognizes that achieving an ideal gait input—one that is both rich in discriminative features and robust to irrelevant variations—remains an open challenge in the field.
$\bullet$ Real-world Challenges. Despite significant data collection efforts, current gait recognition datasets still struggle to comprehensively capture real-world noise factors that influence walking patterns. On one hand, constrained datasets [10, 11, 21, 20, 81, 23] allow for flexible annotation but is hard to replicate the complexities of real-world conditions, such as various occlusions, discontinuous videos, and illumination changes. On the other hand, in-the-wild datasets [9, 8] better reflect practical challenges but face difficulties in accurately labeling variations within sequence pairs. To mitigate these issues, recent studies have explored three key strategies: (a) collecting datasets that focus on specific noise factors, such as different types of occlusions [82], (b) leveraging pedestrian attribute classifiers to obtain automatic annotations for in-the-wild walking sequences [8], and (c) pretraining self-supervised gait models on large-scale unlabeled datasets to enhance robustness against noise factors [44]. OpenGait recognizes this as a critical direction in advancing gait recognition research toward better practicality.
$\bullet$ Gait Metric Learning. Despite substantial accuracy gains across various gait benchmarks, the performance of existing models remains insufficient for real-world deployment. Visual analyses of gait feature distributions [83, 84] reinforce the observation that reducing intra-class variance remains a major challenge in gait recognition. Beyond refining gait backbone architectures, recent research has explored two key strategies to address this issue: (a) re-ranking techniques to improve retrieval accuracy by refining similarity relationships [85, 86], and (b) enhanced loss functions tailored for optimizing gait metric learning [87]. OpenGait recognizes gait metric learning as a crucial factor for improving real-world robustness, though it remains an area that has yet to receive sufficient research attention.
$\bullet$ Gait Temporal Modeling. Previous literature widely recognizes two primary approaches to temporal modeling for gait description: set-based[7, 64, 88] and sequence-based[15, 16, 61] methods. In OpenGait, we also explore a set-based variant of DeepGaitV2, where the only modification is replacing ordered frames with shuffled frames during both training and testing. This adjustment leads to a slight accuracy drop of no more than 5% (70.2% vs. 74.4%), aligning with similar findings in related studies [15, 89]. This suggests that holistic temporal dynamics across the entire frame set, rather than just local dependencies between neighboring frames, play a crucial role in existing methods. Two possible explanations emerge: (a) Holistic temporal dynamics may effectively capture gait details due to the periodic nature of gait cycles [7]. (b) More advanced temporal models are still needed for a finer-grained description of gait motion. OpenGait recognizes gait temporal modeling as an ongoing challenge in gait recognition research.
VIII Conclusions